Question 28 (2 points) What is the formal charge on boron in this compound? CHE F FB-NCHE F H -2 0 +2 +1 -1
Answer #1

Any: formal charge beron CH3 F - B - To T CH3 H FC= V-L-1 es 2 atom - where FC = formal charge V = Total number of valence el